OpenStudy (elisaneedshelp):

The longest length of a sail on Jessica's sailboat is 12 sqrt 40 feet long. Which number is equivalent to this length?

OpenStudy (elisaneedshelp):

\[24\sqrt{40}\] \[24\sqrt{10}\] \[48\sqrt{10}\] \[48\sqrt{40}\]

TheSmartOne (thesmartone):

Can we simplify \(\sf\sqrt{40}\) ?

OpenStudy (elisaneedshelp):

yes

OpenStudy (elisaneedshelp):

but i dont know if in this problem

TheSmartOne (thesmartone):

Break up 40 into the squares of numbers such as 4, 9, 16 See if 40 is divisible by any of those so we can simplify it.

TheSmartOne (thesmartone):

Is 40 divisible by 4? Is it divisible by 9? What is the largest square it is divisible by?

OpenStudy (elisaneedshelp):

4 it is divisble by

TheSmartOne (thesmartone):

Correct, so: \(\sf\Large\sqrt{40}=\sqrt{4\times 10}=\sqrt4\times \sqrt{10}\) What is squareroot of 4 equal to?

OpenStudy (elisaneedshelp):

2

TheSmartOne (thesmartone):

So then \(\sqrt{40}=2\sqrt{10}\)

TheSmartOne (thesmartone):

And then: \(\sf 12\sqrt{40} = 12\times 2 \sqrt{10}\) What is 12 x 2= ?

OpenStudy (elisaneedshelp):

24

TheSmartOne (thesmartone):

So your final answer is \(\sf\Huge 24\sqrt{10}\)

OpenStudy (elisaneedshelp):

Thanks man :)

TheSmartOne (thesmartone):

Np :)